Use the data in the table to calculate the Paasche quantity index. Show all work please.

2005 (base year) Price ($) 2005 (base year) Quantity 2015 Price ($) 2015 Quantity
Bread (loaf) 0.77 50 1.98 55
Eggs (dozen) 1.85 26 2.98 20
Milk (gallon) 0.88 102 1.98 130
Answer #1

Solution -

We know that the formula,

Paasche quantity index = Sum P1Q1 / Sum P0Q1

P1Q1 = 1.98 * 55 + 2.98* 20 + 1.98 * 130

= 108.9 + 59.6 + 257.4

= 425.9

P0Q1 = 0.77 * 55 + 1.85*20 + 0.88 * 130

= 42.35 + 37 + 114.4

= 193.75

Paasche quantity index = ​​425.9 /193.75 * 100

= 2.20 * 100

= 220